1.. Which statement is correct about the changes in potential energy and kinetic energy when two N atoms combine to form an N2 molecule?

Group of answer choices

a. The potential energy becomes more negative and the kinetic energy increases.

b. The potential energy becomes more positive and the kinetic energy decreases.

c. The potential energy becomes more positive and the kinetic energy increases.

d. The potential energy becomes more negative and the kinetic energy decreases.

2.. You take solid ammonium chloride and dissolve it in water. The ammonium chloride solution feels cold. Which statement is true about the process of ammonium chloride dissolving in water?

Group of answer choices

a. In an endothermic reaction, the kinetic energy of the products is lower than the kinetic energy of the reactants. This is why the temperature of the solution drops.

b. The reactants of an endothermic reaction have a higher potential energy than the products. This is why the temperature of the solution drops.

c. An endothermic reaction transfers heat to the surrounding. This is why the temperature of the solution drops.

3..

Use bond energies to determine the energy transferred in the reaction between hydrogen and chlorine to make hydrogen chloride.

H2 (g) + Cl2 (g) --> 2HCl (g)

Bond

Average bond energy (kJ/mol)

H-H

436

Cl-Cl

243

H-Cl

431

Group of answer choices

a. −248 kJ

b. +248 kJ

c. +183 kJ

d. −183 kJ
Answer #1

1. d ) Since the N2 molecule is a very stable molecule. So, the formation of N2 will lead to a more decrease in kinetic energy and potential energy decreases in the potential energy surface diagram.

2. c) b. The reactants of an endothermic reaction have higher potential energy than the products. This is why the temperature of the solution drops The heat supplied will break the NH4Cl and this heat is dissipated to the surroundings.

3. This question is solved by the addition of energy of all the reactants. Then the addition of bond energies in case of products.

As we know bond energy is simply the energy required to break a chemical bond.

So, we need the calculate the difference between the bond energies of reactants and products

Energy of reactants =436 + 243 = 679

Energy of reactants = 2 * 431 = 862

Now difference = 679 - 862 = - 183kj/mol